OpenStudy (dcolley):

Can someone help please Rewrite sec(65°) in terms of its cofunction

OpenStudy (jamesj):

What is the cofunction of sec and what is its relation to sec? Hint: what is the relation between sin and cosin?

OpenStudy (dcolley):

Secant goes with cosign right? Its equal to the cofunction of angles complement

OpenStudy (jamesj):

cosec, yes. Now sin(90-x) = cos(x). Hence 1/cos(x) = 1/sin(90-x) which means by definition of sec and cosec that sec(x) = cosec(90-x)

OpenStudy (jamesj):

Got your answer then? sec(65) = cosec( .....what? ....)